How many milliliters of 0.100 M HCIO3 are required to neutralize 40.0 mL of 0.140 M KOH?
Answer #1

mi ng balanced equation is - kolt & Helog kelog + Hei се клоо жаt H, 4 - 1,4 Miz Molarity of Huag=0-100 Ň Vi= volume of Hulog